# Capacity Gains in MIMO Systems with Few-Bit ADCs Using Nonlinear Analog Circuits

### v2

@@ VII-C Achievable Inner-Bound for Hybrid Beamforming with Envelope Detectors
Figure 14: (a) The circuit design for the generation of fourth and second order polynomials, (b) the power consumption breakdown of the circuits for generation of equal voltage amplitude (corresponding to 0 dBm power) at the second and fourth harmonics.
VIII Circuit Design for Polynomial Operators and Envelope Detectors
In this section, we provide power simulation and measurement results using 22 nm FD SOI CMOS technology and 65nm Bulk CMOS technology. These measurements build upon [55]. Figure 12 shows the architecture of a conventional hybrid beamforming receiver (left) and the proposed receiver (right). It can be noted that the main difference between the two architectures is that the linear analog combiner, implemented in baseband, is replaced by a set of non-linear analog operators.
VIII-A Analysis Based on 65 nm Bulk CMOS Technology
In the following, we provide two circuit designs for envelop detectors and polynomial operators to show the feasibility of the proposed architecture using the 65 nm Bulk CMOS Technology. There are two main advantages to this technology compared to the 22nm technology considered in the sequel: 1) 65 nm technology has a higher nominal supply voltage than 22 nm and if the realized analog operators remain power-efficient, deploying the proposed receiver architecture across a larger group of CMOS technologies is justified. 2) SOI technologies exhibit more threshold voltage variation compared to the bulk CMOS technologies which will impact the operation of baseband circuitry [56]. Meanwhile, the 22 nm technology offers faster transistors which are better choices for the mm-wave signal chain. Therefore, both single-die designs where baseband and mm-wave chain co-exist on the same chip as well as two-die solutions deploying Bulk CMOS for baseband and SOI CMOS for the mm-wave chain can be pursued.
Figure 15: (a) The cascade of circuits emulating absolute value operators, (b) conventional pipeline ADC architecture.
VIII-B Polynomial Function Circuit Design
We assume that we are given a direct current (DC) signal (proportional to the bit value) and our objective is to produce a polynomial function of degree up to four of the input DC value. In practice, there are two methods to achieve this objective: (i) DC domain nonlinear function synthesis based on the quadratic I-V characteristic of the transistor and increasing the order of polynomial by cascading circuits [57], (ii) translating DC values to sinusoidal waveforms, and then generating harmonics of these waveforms whose amplitude is polynomially dependent on the fundamental frequency amplitude. The former has a simpler circuitry; however it can only be used to produce a specific set of polynomials, i.e., restricted polynomial coefficient values. The latter can produce polynomials with arbitrary coefficients through efficient filtering of undesired harmonic terms. However, it comes with higher power consumption and more complex circuitry, and its implementation requires careful characterization of transistors’ nonlinearity, e.g., using Volterra-Weiner series representation methods [58, 59].
@@ VIII-B Polynomial Function Circuit Design
Figure 16: (a) The deployed differential to single-ended op-amp deployed in the envelope detector (b) the circuit diagram of the implemented envelope detector
Figure 17: The impact of amplifier gain on the half-cycle rectification amplitude distortion illustrated for (a) gain of 10 dB, and (b) 25 dB.
To explain the proposed construction, let us consider the problem of producing the fourth order polynomial $f(x)=x^{4}+x^{2}$, where $x$ is the DC input value. Since naturally the amplitude level of the fourth harmonic is less than that of the second harmonic, for a sinusoidal waveform, a harmonic-centric power enhancement at the fourth harmonic is needed to produce the desired polynomial. Fig. 14(a) shows a circuit to generate $f(x)=x^{4}+x^{2}$. In order to generate equal amplitudes at the second and fourth harmonics, the power gain of the transistors generating the fourth harmonic should be larger, leading to an increased power consumption in generating the fourth order term compared to the second order term. Figure 14(b) illustrates numerical values for the power consumption of the proposed circuit through simulations. It can be observed that the ratio of the power consumption for the generation of fourth order term compared to the second order term increases with frequency since the transistor power gain drops at higher frequencies. It is noteworthy that by increasing the baseband bandwidth to support higher data rates, a higher frequency for the generation of polynomials in Fig. 14 may be needed. The simulated results in Fig. 14 are based on 65nm Bulk CMOS technology. The power consumption can be reduced by transitioning into smaller transistor nodes.
Figure 18: The input waveform eye diagram for (a) PAM4 and (b) PAM8 modulations compared with the corresponding envelope detector output eye diagrams in (c) and (d).
VIII-C Implementation of Envelope Detectors
The circuit block diagram of the proposed multi-step envelope detector is shown in Fig. 15(a). Compared with the conventional pipe-line ADC shown in Fig. 15(b) this circuit exhibits major power saving by removing the one-bit DAC and subtractor used in each stage. In terms of functionality, the linearity in both scenarios is mainly limited by gain and bandwidth of operational amplifiers (Op-amp) used in functional blocks.
@@ VIII-C Implementation of Envelope Detectors
Figure 19: The linear growth of power consumption with the data rate in the nonlinear analog operator circuits, i.e., the polynomial generators and envelope detectors.
By cascading multiple stages of the envelope detectors, as mentioned in Section IV the achievable data-rate increases. To justify this claim, we have simulated the two-stage envelope detector circuit in Fig. 15(a). The DC level shifter is realized by diode-based circuits which consume no DC power and can operate up to 6 GHz.
The simulated eye-diagram performance of the two-stage envelope detector is illustrated in Fig. 18. It can be noted that at such higher data rates (12 Gb/s), in contrast to lower data rates where a square wave is feasible, sinusoidal waveforms are the only feasible inputs to an ADC [60]. Therefore, in our simulations we consider the input to the absolute value function circuitry to be a sinusoidal waveform due to the slewing effects [61].
An important characteristic of the envelope detectors is the input dynamic range to support an output waveform that follows the input amplitude with minimal deviation. By increasing the input magnitude, the amplifier’s constituent transistors will be pushed into nonlinear regions of operation, and a “compression behavior” is observed. The point at which this compression happens is an upper-bound on the magnitude of input signals. This value is critical in scenarios where amplitude modulation is deployed in the transceiver. To evaluate our proposed envelope detector, two modulation scenarios of Pulse Amplitude Modulation (PAM)4 and PAM8 modulation schemes are considered as the input to the envelope detector, as shown in Fig. 18. In both cases, the output rectified eye diagram exhibits clear distinction between the amplitude levels (denoted by different colors) such that the comparators in the following one-bit ADCs can distinguish the amplitude levels with small error. The simulation results in Fig. 18 demonstrate that the amplitude ratios at the output follow the input amplitude ratios even when the dynamic range of input waveform is only between -400 to +400 mV. I can be noted that a PAM-8 waveform in the receiver base-band can be constructed by passing the QAM-64 signal (in Table I) through a quadrature down-conversion mixer [62].
Theorems 2 and 4 show that the channel capacity depends on the number of ADCs through $n_{q}2^{\delta_{env}}+1$ and $n_{q}\delta_{poly}+1$, respectively, so that the use of a envelope detectors and quadratic analog operators instead of a linear operators ($\delta_{env}:0\to 1$ and $\delta_{poly}:1\to 2$) has an equivalent effect on capacity as that of quadrupling and doubling the number of ADCs $n_{q}$, respectively. This fact, along with the power consumption values given in Figures 13 and 19 justify the use of nonlinear analog operators. It is understood that power consumption is dependent on circuit configuration, transistor size, and passive quality factors. However, the values reported in this paper based on recent demonstrations of a mm-wave transceiver in a 22 nm FDSOI CMOS technology serve as a proof-of-concept to justify the effectiveness of the proposed receiver architecture designs.
VIII-D Analysis Based on 22 nm FDSOI CMOS Technology
In the Figure 13, the itemized and total power consumption is compared between the two architectures in a 22 nm FDSOI CMOS technology. For the power values in Figur 13, we assume a mm-wave carrier frequency of 60 GHz, RF bandwidth of 1 GHz, and 16 channels. The reported power consumption for RF elements are based on the measurement results of the chip reported in [55] and the power consumption of nonlinear operators and ADCs are based on post-layout simulation results and analytical values in [63], respectively.
@@ IX Conclusion
The application of nonlinear analog operations in MIMO receivers was considered. A receiver architecture consisting of linear analog combiners, implementable nonlinear analog operators, and few-bit threshold ADCs was designed, and the fundamental information theoretic performance limits of the resulting communication system were investigated. Furthermore, circuit-level simulations, using a 22 nm FDSOI CMOS technology, were provided to show the implementability of the desired nonlinear analog operators with practical power budgets.
@@ Appendix A Proof of Proposition 1
Let $n_{q}=\lfloor\frac{P_{ADC}}{2\alpha}\rfloor$. For a given quantizer with associated code $\mathcal{C}$, the high SNR achievable rate is $\log{|\mathsf{P}|}=\log{|\mathcal{C}|}$. So, finding the capacity is equivalent to finding the maximum $|\mathcal{C}|$ over all choices of $Q(\cdot)$. First, let us prove the converse result. Note that $|\mathcal{C}|\leq 2n_{q}$ since $\mathbf{c}_{0}=\mathbf{c}_{2n_{q}}$. The reason is that for the absolute value function $f_{a_{j},j}(\cdot),j\in[n_{q}]$, we have $\lim_{y\to\infty}f_{a_{j},j}(y)=\lim_{y\to-\infty}f_{a_{j},j}(y)=\infty$. So,
$c_{0,j}=\lim_{y\to-\infty}\mathbbm{1}(f_{a_{j},j}-t_{j}>0)=\lim_{y\to\infty}\mathbbm{1}(f_{a_{j},j}-t_{j}>0)=c_{2n_{q},j}.$ (20)
As a result, $\log{|\mathcal{C}|}\leq 1+\log{n_{q}}$. Next, we prove achievability. Let $t_{j}=\frac{n_{q}+1}{2},j\in[n_{q}]$ and $f_{a_{j},j}(y)\triangleq|y-j-\frac{n_{q}+1}{2}|,j\in[n_{q}]$, so that the roots of $f_{a_{j},j}(\cdot)-t_{j}$ are $j$ and $j-n_{q}-1$. Then, $(r_{1},r_{2},\cdots,r_{2n_{q}})\!=\!(-n_{q},-n_{q}\!+\!1,\cdots,-1,1,2,\cdots,n_{q})$ and
$c(i,j)=\begin{cases}1-\mathbbm{1}(j\leq i)\qquad&\text{if }\quad i\leq n_{q},\\ \mathbbm{1}(n_{q}-j+1\leq i-n_{q})&\text{otherwise}.\end{cases}$
@@ Appendix B Proof of Theorem 1
We provide an outline of the proof. Note that the input alphabet has at most $12n_{q}+$ 1 mass points since based on the proof of Proposition 1, the channel output can take at most $2n_{q}$ values. Let the quantized channel output be denoted by $\widehat{Y}$. Since the conditional measure $P_{\widehat{Y}|X}(\cdot|x),x\in\mathbb{R}$ is continuous in $x$, and $\lim_{x\to\infty}P_{\widehat{Y}|X}(\mathcal{A}|x)=\mathbbm{1}(\hat{y}\in\mathcal{A}),\mathcal{A}\in\mathbb{B}$ for some fixed $\hat{y}$, the conditions in the proof of [64, Prop. 1] hold, and the optimal input distribution has bounded support. From the extension of Witsenhausen’s result [65] given in [64, Prop. 2], the optimal input distribution is discrete and takes at most ${2n_{q}+1}$ values. This completes the proof of converse. To prove achievability, it suffices to show that one can choose the set of functions $f_{a_{j},j}(\cdot),j\in[n_{q}]$ and quantization thresholds $t_{j},j\in[n_{q}]$ such that the resulting quantizer operates as described in the theorem statement, i.e., it generates $\widehat{Y}=Q(hX+N)$ where $Q(y)=k$ if $y\in[t_{k},t_{k+1}],k\in\{1,\cdots,2n_{q}]$ and $Q(y)=0$ if $y>t_{2n_{q}}$ or $y<t_{1}$. To this end, let $\mathbf{t}^{*}$ be the optimal quantizer thresholds in (2). Let $r_{1},r_{2},\cdots,r_{2n_{q}}$ be the elements of $\mathbf{t}^{*}$ written in non-decreasing order. Define a quantizer with associated analog functions $f_{a,j}(y)\triangleq|y-\frac{r_{j}+r_{n_{q}+j}}{2}|$ and $t_{j}=\frac{r_{n_{q}+j}-r_{j}}{2}$. Note that $t_{j}>0$ since $r_{j},j\in[n_{q}]$ are non-decreasing. Then, similar to the proof of Proposition 1, the quantization rule gives distinct outputs for $y\in[r_{k},r_{k+1}],k\in\{1,\cdots,2n_{q}]$ and $y\in[r_{2n_{q}},\infty)\cup[-\infty,r_{1}]$ as desired. ∎
@@ Appendix C Proof of Proposition 5
where $\eta_{s}\triangleq 2^{\delta}-\sum_{s^{\prime}=1}^{s-1}2^{\delta-s^{\prime}}+1,s>1$. Consider a quantizer $Q(\cdot)$ with ADC thresholds $t(1:n_{q})$ and associated analog functions $f_{j}(y)\triangleq A_{\delta}(x,a^{\delta}),j\in[n_{q}]$. Then, $r_{1},r_{2},\cdots,r_{\gamma}$ are the non-decreasing sequence of roots of $f_{j}(\cdot),j\in[n_{q}]$, and the associated code of the quantizer $Q(\cdot)$ is $\mathcal{C}$ as desired. ∎

### v3

@@ VIII Circuit Design and Power Analysis
Figure 9: (a) Generation of fourth and second order polynomials. (b) Power consumption for generation of equal voltage amplitude (0 dBm power) at the second and fourth harmonics. (c) The cascade of circuits emulating absolute value operators.
@@
Figure 10: (a) Differential to single-ended op-amp in the envelope detector (b) Circuit diagram of the envelope detector
Figure 7 shows the architecture of a conventional hybrid beamforming receiver (left) and the proposed receiver (right). In the Figure 8, the itemized and total power consumption is compared between the two architectures in a 22 nm FDSOI CMOS technology. For the power values in Figure 8, we assume a mm-wave carrier frequency of 60 GHz, RF bandwidth of 1 GHz, and 16 channels. The reported power consumption for RF elements are based on the measurement results of the chip reported in [41] and the power consumption of nonlinear operators and ADCs are based on post-layout simulation results and analytical values in [42], respectively. To simplify the rate analysis we assume communication takes place at the high SNR regime (Theorems 3, 4, and 5).
VIII-A Analysis Based on 22 nm FDSOI CMOS Technology
@@
[…]
Figure 11: (a) Impact of amplifier gain on the half-cycle rectification amplitude distortion for gain of 10 dB (left), and 25 dB (right). (b) Input waveform eye diagram for (top left) PAM4 and (bottom left) PAM8 modulations compared with that of setups with envelop detectors (right).
VIII-B Analysis Based on 65 nm Bulk CMOS Technology
We provide two circuit designs for envelope detectors and polynomial operators using the 65 nm Bulk CMOS technology. There are two main advantages to this technology: 1) 65 nm technology has a higher nominal supply voltage than 22 nm. 2) SOI technologies exhibit more threshold voltage variation compared to the bulk CMOS technologies which impacts the operation of baseband circuitry [43].
Polynomial Operators: To explain the proposed construction, consider $f(x)=x^{4}+x^{2}$, where $x$ is the input value. Fig. 9(a) shows a circuit to generate $f(x)=x^{4}+x^{2}$. In order to generate equal amplitudes at the second and fourth harmonics, the power gain of the transistors generating the fourth harmonic should be larger, leading to an increased power consumption in generating the fourth order term compared to the second order term. Figure 9(b) illustrates numerical values for the power consumption of the proposed circuit through simulations. For large bandwidth communications, a higher frequency for the generation of polynomials in Figure 9(a) may be needed.
[…]
Simulation of two concatenated envelope detector circuits is shown in Figure 9(c). The DC level shifter is realized by diode-based circuits which consume no DC power and can operate up to 6 GHz. The simulated eye-diagram performance of the two-stage envelope detector is illustrated in Figure 11(b). For high data rates (12 Gb/s), we consider the input to be the absolute value function circuitry to be a sinusoidal waveform due to the slewing effects [44]. To evaluate our proposed envelope detector, two modulation scenarios of Pulse Amplitude Modulation PAM4 and PAM8 modulation schemes are considered as the input to the envelope detector, as shown in Figure 11(b). The simulation results demonstrate that the amplitude ratios at the output follow the input amplitude ratios even when the dynamic range of input waveform is only between -400 to +400 mV.
@@ IX Conclusion
The application of nonlinear analog operations in MIMO receivers was considered. A receiver architecture consisting of linear analog combiners, implementable nonlinear analog operators, and few-bit threshold ADCs was designed, and the fundamental information theoretic performance limits of the resulting communication system were investigated. Furthermore, circuit-level simulations, were provided to show the implementability of the desired nonlinear analog operators with practical power budgets.
@@ Appendix A Proof of Proposition 1
Let $n_{q}=\lfloor\frac{P_{ADC}}{2\alpha}\rfloor$. For a given quantizer with associated code $\mathcal{C}$, the high SNR achievable rate is $\log{|\mathsf{P}|}=\log{|\mathcal{C}|}$. So, finding the capacity is equivalent to finding the maximum $|\mathcal{C}|$ over all choices of $\mathbf{q}(\cdot)$. First, let us prove the converse result. Note that $|\mathcal{C}|\leq 2n_{q}$ since $\mathbf{c}_{0}=\mathbf{c}_{2n_{q}}$. The reason is that for the absolute value function $f_{j,1}(\cdot),j\in[n_{q}]$, we have $\lim_{y\to\infty}f_{j,1}(y)=\lim_{y\to-\infty}f_{j,1}(y)=\infty$. So,
$c_{1,j}=\lim_{y\to-\infty}\mathbbm{1}(f_{j}-t_{j,1}>0)=\lim_{y\to\infty}\mathbbm{1}(f_{j}-t_{j,1}>0)=c_{2n_{q},j}.$ (17)
As a result, $\log{|\mathcal{C}|}\leq 1+\log{n_{q}}$. Next, we prove achievability. Let $t_{j,1}=\frac{n_{q}+1}{2},j\in[n_{q}]$ and $f_{j,1}(y)\triangleq|y-j-\frac{n_{q}+1}{2}|,j\in[n_{q}]$, so that the roots of $f_{j}(\cdot)-t_{j,1}$ are $j$ and $j-n_{q}-1$. Then, $(r_{1},r_{2},\cdots,r_{2n_{q}})\!=\!(-n_{q},-n_{q}\!+\!1,\cdots,-1,1,2,\cdots,n_{q})$ and
$c_{i,j}=\begin{cases}1-\mathbbm{1}(j\leq i)\qquad&\text{if }\quad i\leq n_{q},\\ \mathbbm{1}(n_{q}-j+1\leq i-n_{q})&\text{otherwise}.\end{cases}$
@@ Appendix B Proof of Theorem 1
We provide an outline of the proof. Note that the input alphabet has at most $12n_{q}+$ 1 mass points since, based on the proof of Proposition 1, the channel output can take at most $2n_{q}$ values. Let the quantized channel output be denoted by $\widehat{W}$. Since the conditional measure $P_{\widehat{W}|X}(\cdot|x),x\in\mathbb{R}$ is continuous in $x$, and $\lim_{x\to\infty}P_{\widehat{W}|X}(\mathcal{A}|x)=\mathbbm{1}(\hat{w}\in\mathcal{A}),\mathcal{A}\in\mathbb{B}$ for some fixed $\hat{w}$, the conditions in the proof of [45, Prop. 1] hold, and the optimal input distribution has bounded support. From the extension of Witsenhausen’s result [46] given in [45, Prop. 2], the optimal input distribution is discrete and takes at most ${2n_{q}+1}$ values. This completes the proof of converse. To prove achievability, it suffices to show that one can choose the set of functions $f_{j}(\cdot),j\in[n_{q}]$ and quantization thresholds $t_{j,1},j\in[n_{q}]$ such that the resulting quantizer operates as described in the theorem statement, i.e., it generates $\widehat{W}=\mathbf{q}(hX+N)$ where $\mathbf{q}(y)=k$ if $y\in[t_{k},t_{k+1}],k\in\{1,\cdots,2n_{q}]$ and $\mathbf{q}(y)=0$ if $y>t_{2n_{q}}$ or $y<t_{1}$. To this end, let $\mathbf{t}^{*}$ be the optimal quantizer thresholds in (9). Let $r_{1},r_{2},\cdots,r_{2n_{q}}$ be the elements of $\mathbf{t}^{*}$ written in non-decreasing order. Define a quantizer with associated analog functions $f_{j,1}(y)\triangleq|y-\frac{r_{j}+r_{n_{q}+j}}{2}|$ and $t_{j,1}=\frac{r_{n_{q}+j}-r_{j}}{2}$. Note that $t_{j}>0$ since $r_{j},j\in[n_{q}]$ are non-decreasing. Then, similar to the proof of Proposition 1, the quantization rule gives distinct outputs for $y\in[r_{k},r_{k+1}],k\in\{1,\cdots,2n_{q}]$ and $y\in[r_{2n_{q}},\infty)\cup[-\infty,r_{1}]$ as desired. ∎
@@ Appendix C Proof of Proposition 5
where $\eta_{s}\triangleq 2^{\delta}-\sum_{s^{\prime}=1}^{s-1}2^{\delta-s^{\prime}}+1,s>1$. Consider a quantizer $\mathbf{q}(\cdot)$ with ADC thresholds $t(1:n_{q})$ and associated analog functions $f_{j}(y)\triangleq A_{\delta}(x,a^{\delta}),j\in[n_{q}]$. Then, $r_{1},r_{2},\cdots,r_{\gamma}$ are the non-decreasing sequence of roots of $f_{j}(\cdot),j\in[n_{q}]$, and the associated code of the quantizer $\mathbf{q}(\cdot)$ is $\mathcal{C}$ as desired. ∎